OpenStudy (anonymous):

Which statement shows a correct next step in solving the equation? The equation can become 6x + 1 = 1 by combining like terms. please help?

OpenStudy (mathstudent55):

I think a lot of info is missing.

OpenStudy (anonymous):

yes im sorry is An equation is shown below: (4x - 1) + (2x + 2) = 1 Which statement shows a correct next step in solving the equation? The equation can become 6x + 1 = 1 by combining like terms. .The equation can become 3x + 4x = 1 by solving the terms in the brackets. . The equation can become 3x + 4x = 1 by applying the associative property of multiplication. The equation can become 6x + 1 = 1 by applying the distributive property.

OpenStudy (anonymous):

@mathstudent55